# Recursive Joint Cross-Modal Attention for Multimodal Fusion in Dimensional Emotion Recognition

### v2

@@ Abstract
Though multimodal emotion recognition has achieved significant progress over recent years, the potential of rich synergic relationships across the modalities is not fully exploited. In this paper, we introduce Recursive Joint Cross-Modal Attention (RJCMA) to effectively capture both intra- and inter-modal relationships across audio, visual and text modalities for dimensional emotion recognition. In particular, we compute the attention weights based on cross-correlation between the joint audio-visual-text feature representations and the feature representations of individual modalities to simultaneously capture intra- and inter-modal relationships across the modalities. The attended features of the individual modalities are again fed as input to the fusion model in a recursive mechanism to obtain more refined feature representations. We have also explored Temporal Convolutional Networks (TCNs) to improve the temporal modeling of the feature representations of individual modalities. Extensive experiments are conducted to evaluate the performance of the proposed fusion model on the challenging Affwild2 dataset. By effectively capturing the synergic intra- and inter-modal relationships across audio, visual and text modalities, the proposed fusion model achieves a Concordance Correlation Coefficient (CCC) of 0.585 (0.542) and 0.659 (0.619) for valence and arousal respectively on the validation set (test set). This shows a significant improvement over the baseline of 0.24 (0.211) and 0.20 (0.191) for valence and arousal respectively on the validation set (test set) of the valence-arousal challenge of 6th Affective Behavior Analysis in-the-Wild (ABAW) competition. The code is available on GitHub: https://anonymous.4open.science/r/RJCMA-89A5.
@@ 1 Introduction
Emotion recognition is a challenging problem due to the diverse nature of expressions associated with emotional states across individuals and cultures [1]. It has a wide range of applications in various fields such as health care (assessing pain, fatigue, depression) [2, 3], autonomous driving (assessment of driver emotional states) [4], robotics (for realistic human-computer interaction) [5], etc. Emotion recognition is often explored in the literature as a classification problem of categorizing the emotions to seven classes: anger, disgust, fear, happy, sad, surprise and contempt [6]. Recently, compound expression datasets have also been introduced to capture human emotions beyond seven classes [7]. However, they also do not capture the diverse range of emotions expressed by humans. Therefore, dimensional model of emotions has been introduced to capture the diverse range of emotions, predominantly using valence and arousal as shown in Fig 1. Valence denotes the range of emotions from very sad (negative) to very happy (positive), whereas arousal represents the intensities of emotions from being very passive (sleepiness) to very active (high excitement) [8]. Dimensional Emotion Recognition (DER) is more challenging than categorical emotion recognition, as they are highly prone to label noise due to complex process of obtaining dimensional annotations.
Figure 1: The valence-arousal space [9]
@@
Multimodal learning has been recently gaining a lot of attention as they can offer rich complementary information across multiple modalities, which can play a crucial role to outperform unimodal approaches [10]. Human emotions are communicated in complex ways through various modalities such as face, voice, and language. Multimodal emotion recognition aims to effectively fuse the audio, visual and text by capturing the rich intra- and inter-modal complementary relationships across the modalities. Early approaches of multimodal fusion for DER either rely on LSTM-based fusion [11, 12] or early feature concatenation [13, 14]. With the advent of transformers [15], attention models using multimodal transformers have attained much interest to effectively combine multiple modalities for DER [16, 17, 18]. Recently, cross-modal attention was found to be quite promising in capturing the rich inter-modal complementary relationships across the modalities, which has been successfully applied for several applications such as action localization [19], emotion recognition [20], and person verification [21]. Praveen et al [9, 22] explored joint cross-attention by introducing joint feature representation in the cross-attentional framework and achieved significant improvement over vanilla cross-attention [20]. They have further improved the performance of the model by introducing recursive fusion, and LSTMs for temporal modeling of individual feature representations as well as the audio-visual representations [23]. Recursive fusion of the cross-attention models has also been successfully explored for other audio-visual tasks such as event localization [24] and person verification [25]. However, most of the existing cross-attention based models are focused on audio-visual fusion for DER.
@@
In this work, we have investigated the prospect of efficiently capturing the synergic intra- and inter-modal relationships across audio, visual, and text modalities for DER. By deploying cross-correlation between the joint audio-visual-text feature representation and feature representations of individual modalities, we are able to simultaneously capture intra- and inter-modal relationships across the modalities. Inspired by the performance of recursive attention models [24, 23], we have also incorporated recursive fusion of audio, visual and text modalities in the context of joint cross-attentional fusion to obtain more refined feature representations. The major contributions of the paper can be summarized as follows.
@@
  • Joint cross-modal attention is explored among audio, visual and text modalities using a joint audio-visual-text feature representation to simultaneously capture intra- and inter-modal relationships across the modalities.
@@
  • Extensive experiments are conducted to evaluate the robustness of the proposed approach in the challenging Affwild2 dataset.
@@ 2.1 Multimodal Emotion Recognition
One of the early approaches using deep learning architectures for DER was proposed by Tzirakis et al. [11], where they explored 1D Convolutional Neural Networks (CNN) for audio and Resnet-50 [26] for visual modality. The deep features are then fused using Long Short Term Memory (LSTM) Networks for estimating predictions of valence and arousal. With the advancement of 3D CNN models, Kuhnke et al. [13] showed performance improvement using R3D [27] for visual modality and Resnet-18 for audio modality with simple feature concatenation. Kollias et al. [28, 29, 30] explored DER along with other tasks of classification of expressions and action units in a unified framework. Another widely explored line of research for DER is based on knowledge distillation (KD) [31]. Schoneveld et al. [12] explored KD for obtaining robust visual representations, while Wang et al. [32] and Deng et al. [33] attempted to leverage KD to deal with label uncertainties. Recently, KD is also explored along with the paradigm of Learning Under Privileged Information (LUPI) to efficiently exploit multiple modalities for DER [34]. Li et al. [35] proposed Decoupled Multimodal Distillation (DMD) to mitigate the issues of multimodal heterogeneities by dynamically distilling the modality relevant information across the modalities. Although these methods have shown promising performance by exploiting multiple modalities, they do not focus on capturing the synergic relationships pertaining to intra- and inter-modal relationships across the modalities.
@@ 2.2 Attention Models for Multimodal Emotion Recognition
Inspired by the performance of transformers [15], several approaches have been proposed to investigate the potential of transformers for DER. Most of the existing works explored transformers to encode the concatenated version of the feature representations of individual modalities [17, 18, 36, 37]. Tran et al. [38] showed that fine-tuning transformers with cross-modal attention trained on a large-scale voxceleb2 dataset helps to improve the performance of multimodal emotion recognition. Huang et al. [39] explored multimodal transformers along with self-attention modules, where audio modality is used to attend to visual modality to produce robust multimodal feature representations. Parthasarathy et al. [16] further extended their idea by employing the cross-modal attention in a bidirectional fashion, where audio modality is used to attend to visual modality and vice-versa. Karas et al [40] provided a comprehensive evaluation of the fusion models based on self-attention, cross-attention and LSTMs for DER. Zhang et al [41] proposed leader-follower attention, where audio, visual and text modalities are combined to obtain the attended feature using the modality-wise attention scores, which is further concatenated to visual features for final predictions.
@@ 3.1 Visual Network
Facial expressions plays a significant role in conveying the emotional state of a person. In videos, spatial information provide the semantic regions of face pertinent to the expressions, whereas temporal dynamics convey the evolution of expressions in videos across the frames. Therefore, effectively modeling the spatial and temporal dynamics of facial expressions in videos is crucial to obtain robust visual feature representations. Several approaches have been explored using 2D CNNs in conjunction with LSTMs, where 2D CNNs are used to encode the spatial information and LSTMs for temporal dynamics of facial expressions [44, 45]. With the advent of 3D CNN models [27], they are successfully explored for DER [13, 9, 22] and showed improvement over 2D CNNs with LSTMs. It has also been shown that 3DCNNs in combination with LSTMs are effective in capturing the temporal dynamics [23], where 3DCNNs are efficient in capturing short-term dynamics and LSTMs are robust in modeling long-term dynamics. Recently, TCNs are found to be promising to effectively capture the temporal dynamics for DER [41, 36]. In this work, we have used Resnet-50 [26] pretrained on MS-CELEB-1M dataset [46], which is further fine-tuned on FER+ [42] dataset similar to that of [41]. we have further used TCNs to effectively capture the temporal dynamics of facial expressions.
@@ 3.2 Audio Network
Speech-based emotion recognition is another promising research area due to the rich emotion-relevant information in the vocal expressions. With the advancement of deep learning models, vocal expressions are encoded using 1D CNNs with raw speech signals [11] or 2D CNNs with spectrograms [12, 13]. Some of the works also explored deep features in combination with conventional hand-crafted features to encode the vocal expressions [17, 47]. Recently, spectrograms have been widely explored as they are found to efficient in capturing the affective states of the vocal expressions [41, 9]. Therefore, we have also explored spectrograms with 2D CNNs to encode the vocal expressions. Specifically, we have used VGG-Net architecture pretrained on large-scale audioset dataset [48]. Similar to visual modality, we have also used TCNs to encode the temporal dynamics of frame-level vocal embeddings.
@@ 3.3 Text Network
Text modality is another predominantly explored modality for emotion detection, which carries semantic emotion-relevant information in the text data [49]. Effectively leveraging the textual data can boost the performance of multimodal fusion as they can offer significant emotion-relevant information and complement audio and visual modalities. Based on transformers, BERT features are predominantly explored text encoders for emotion recognition in the literature [50]. Therefore, we also used BERT as text encoder, followed by TCNs to encode the temporal information across the word embeddings similar to that of [41].
@@ 3.4 Recursive Joint Cross-Modal Attention
Given the input video sub-sequence $S$ of $K$ frames, the audio, visual and text data are preprocessed and fed to the corresponding encoders, followed by TCNs to obtain the feature representations of the respective modalities as shown in Fig 2. The feature representations of the audio, visual and text modalities are denoted by ${\boldsymbol{X}}_{a}\boldsymbol{=}\{\boldsymbol{x}_{a}^{1},\boldsymbol{x}_{a}^{2},...,\boldsymbol{x}_{a}^{K}\boldsymbol{\}}\in\mathbb{R}^{d_{a}\times K}$, ${\boldsymbol{X}}_{v}\boldsymbol{=}\{\boldsymbol{x}_{v}^{1},\boldsymbol{x}_{v}^{2},...,\boldsymbol{x}_{v}^{K}\boldsymbol{\}}\in\mathbb{R}^{d_{v}\times K}$ and ${\boldsymbol{X}}_{t}\boldsymbol{=}\{\boldsymbol{x}_{t}^{1},\boldsymbol{x}_{t}^{2},...,\boldsymbol{x}_{t}^{K}\boldsymbol{\}}\in\mathbb{R}^{d_{t}\times K}$ respectively, where $d_{a}$, $d_{v}$, and $d_{t}$ are dimensions of audio, visual and text features. $\boldsymbol{x}_{a}^{k}$, $\boldsymbol{x}_{v}^{k}$ and $\boldsymbol{x}_{t}^{k}$ represents the feature vectors of the individual frames of audio, visual and text modalities.
@@ 4.1 Dataset
Affwild is the largest database in the field of affective computing, consisting of $298$ videos captured under extremely challenging conditions from YouTube videos [51]. The database is extended substantially to foster the development of robust models for estimating valence and arousal by introducing $260$ more videos, resulting in total of $558$ videos with $1.4$M frames [52]. The database has been further expanded by adding new videos in the subsequent series of ABAW challenges [53, 54, 55, 56, 57, 58]. In the 6th ABAW challenge [59] for the valence-arousal track, the dataset is provided with $594$ videos of around $2,993,081$ frames from $584$ subjects. Sixteen of these videos display two subjects, both of which have been annotated. The final annotations are obtained by taking an average of annotations provided by four experts, using a joystick. The annotations for valence and arousal are provided continuously in the range of $[-1,1]$. The dataset is partitioned into the training, validation and test sets in a subject-independent manner to ensure that each subject appears exclusively in one partition. The partitioning resulted in $356$, $76$, and $162$ videos for train, validation and test partitions respectively.
@@ 4.2.1 Preprocessing
For the visual modality, we have used the cropped and aligned images provided by the challenge organizers [59]. For the missing faces in the video frames, we have considered black frames (i.e., zero pixels) and the video frames with no annotations of valence and arousal i.e., with annotations of $-5$ are excluded. The given video sequences are divided into sub-sequences of $300$ (i.e.,$K=300$) with a stride of $200$ frames and the facial images are resized to $48\times 48$.
@@
For the text modality, the extracted speech signals from audio preprocessing are fed to the pretrained speech recognition model of Vosk toolkit [^2] https://alphacephei.com/vosk/models/vosk-model-en-us-0.22.zip to obtain the recognized words along with word-level timestamps. Next, a pretrained punctuation restoration and capitalization model [^3] https://pypi.org/project/deepmultilingualpunctuation/ is used to restore the punctuations of the recognized words, which carries semantic information pertinent to emotional states. Now BERT features are extracted at word-level using a pretrained BERT model [^4] https://pypi.org/project/pytorch-pretrained-bert/. The word-level features are computed by taking a summation of last four layers of the BERT model [60]. The recognized words may usually span a larger time-window of multiple frames. In order to synchronize the word-level BERT features of text modality with other audio and visual modalities, the word-level text embedding is populated as per the timestamp of each word by reassigning the same word-level feature to all the frames within the time span of the corresponding word.

### v4

@@ Abstract
Though multimodal emotion recognition has achieved significant progress over recent years, the potential of rich synergic relationships across the modalities is not fully exploited. In this paper, we introduce Recursive Joint Cross-Modal Attention (RJCMA) to effectively capture both intra- and inter-modal relationships across audio, visual, and text modalities for dimensional emotion recognition. In particular, we compute the attention weights based on cross-correlation between the joint audio-visual-text feature representations and the feature representations of individual modalities to simultaneously capture intra- and inter-modal relationships across the modalities. The attended features of the individual modalities are again fed as input to the fusion model in a recursive mechanism to obtain more refined feature representations. We have also explored Temporal Convolutional Networks (TCNs) to improve the temporal modeling of the feature representations of individual modalities. Extensive experiments are conducted to evaluate the performance of the proposed fusion model on the challenging Affwild2 dataset. By effectively capturing the synergic intra- and inter-modal relationships across audio, visual, and text modalities, the proposed fusion model achieves a Concordance Correlation Coefficient (CCC) of 0.585 (0.542) and 0.674 (0.619) for valence and arousal respectively on the validation set (test set). This shows a significant improvement over the baseline of 0.240 (0.211) and 0.200 (0.191) for valence and arousal, respectively, in the validation set (test set), achieving second place in the valence-arousal challenge of the 6th Affective Behavior Analysis in-the-Wild (ABAW) competition. The code is available on GitHub: https://github.com/praveena2j/RJCMA.
@@ 1 Introduction
Emotion recognition is a challenging problem due to the diverse nature of expressions associated with emotional states across individuals and cultures [1]. It has a wide range of applications in various fields such as health care (assessing pain, fatigue, depression) [2, 3], autonomous driving (assessing driver emotional states) [4], robotics (for realistic human-computer interaction) [5], etc. Emotion recognition is often explored in the literature as a classification problem of categorizing emotions into seven classes: anger, disgust, fear, happiness, sadness, surprise, and contempt [6]. Recently, compound expression datasets have also been introduced to capture human emotions beyond seven classes [7]. However, they also do not capture the diverse range of emotions expressed by humans. Therefore, a dimensional model of emotions has been introduced to capture the diverse range of emotions, predominantly using valence and arousal. Valence denotes the range of emotions from very sad (negative) to very happy (positive), whereas arousal represents the intensities of emotions from being very passive (sleepiness) to very active (high excitement) [8]. Dimensional Emotion Recognition (DER) is more challenging than categorical emotion recognition, as they are highly prone to label noise due to the complex process of obtaining dimensional annotations.
@@
Multimodal learning has been recently gaining a lot of attention as it can offer rich complementary information across multiple modalities, which can play a crucial role in outperforming unimodal approaches [9]. Human emotions are communicated in complex ways through various modalities such as face, voice, and language. Multimodal emotion recognition aims to effectively fuse audio, visual, and text by capturing the rich intra- and inter-modal complementary relationships across the modalities. Early approaches of multimodal fusion for DER either rely on Long Short Term Memory (LSTM)-based fusion [10, 11] or early feature concatenation [12, 13]. With the advent of transformers [14], attention models using multimodal transformers have attained much interest to effectively combine multiple modalities for DER [15, 16, 17]. Recently, cross-modal attention was found to be quite promising in capturing the rich inter-modal complementary relationships across the modalities, which has been successfully applied for several applications such as action localization [18], emotion recognition [19], and person verification [20]. Praveen et al. [21, 22] explored joint cross-attention by introducing joint feature representation in the cross-attentional framework and achieved significant improvement over vanilla cross-attention [19]. They have further improved the performance of the model by introducing recursive fusion, and LSTMs for temporal modeling of individual feature representations as well as the audio-visual representations [23]. Recursive fusion of the cross-attention models has also been successfully explored for other audio-visual tasks such as event localization [24] and person verification [25]. However, most of the existing cross-attention based models are focused on audio-visual fusion for DER.
@@
In this work, we have investigated the prospect of efficiently capturing the synergic intra- and inter-modal relationships across audio, visual, and text modalities for DER. By deploying cross-correlation between the joint audio-visual-text feature representation and feature representations of individual modalities, we can simultaneously capture intra- and inter-modal relationships across the modalities. Inspired by the performance of recursive attention models [24, 23], we have also incorporated a recursive fusion of audio, visual, and text modalities in the context of joint cross-attentional fusion to obtain more refined feature representations. The major contributions of the paper can be summarized as follows.
@@
  • Joint cross-modal attention is explored among audio, visual, and text modalities using a joint audio-visual-text feature representation to simultaneously capture intra- and inter-modal relationships across the modalities.
@@
  • Extensive experiments are conducted to evaluate the robustness of the proposed approach on the challenging Affwild2 dataset.
@@ 2.1 Multimodal Emotion Recognition
One of the early approaches using deep learning architectures for DER was proposed by Tzirakis et al. [10], where they explored 1D Convolutional Neural Networks (CNN) for audio and Resnet-50 [26] for visual modality. The deep features are then fused using LSTMs for estimating predictions of valence and arousal. With the advancement of 3D CNN models, Kuhnke et al. [12] showed performance improvement using R3D [27] for visual modality and Resnet-18 [26] for audio modality with simple feature concatenation. Kollias et al. [28, 29, 30] explored DER along with other tasks of classification of expressions and action units in a unified framework. Another widely explored line of research for DER is based on knowledge distillation (KD) [31]. Schoneveld et al. [11] explored KD for obtaining robust visual representations, while Wang et al. [32] and Deng et al. [33] attempted to leverage KD to deal with label uncertainties. Recently, KD is also explored along with the paradigm of Learning Under Privileged Information (LUPI) to efficiently exploit multiple modalities for DER [34]. Li et al. [35] proposed Decoupled Multimodal Distillation (DMD) to mitigate the issues of multimodal heterogeneities by dynamically distilling the modality-relevant information across the modalities. Although these methods have shown promising performance by exploiting multiple modalities, they do not focus on capturing the synergic relationships pertaining to intra- and inter-modal relationships across the modalities.
@@ 2.2 Attention Models for Multimodal Emotion Recognition
Inspired by the performance of transformers [14], several approaches have been proposed to investigate the potential of transformers for DER. Most of the existing works explored transformers to encode the concatenated version of the feature representations of individual modalities [16, 17, 36, 37]. Tran et al. [38] showed that fine-tuning transformers with cross-modal attention trained on a large-scale voxceleb2 [39] dataset helps to improve the performance of multimodal emotion recognition. Huang et al. [40] explored multimodal transformers along with self-attention modules, where audio modality is used to attend to visual modality to produce robust multimodal feature representations. Parthasarathy et al. [15] further extended their idea by employing cross-modal attention in a bidirectional fashion, where audio modality is used to attend to visual modality and vice-versa. Karas et al. [41] provided a comprehensive evaluation of the fusion models based on self-attention, cross-attention and LSTMs for DER. Zhang et al. [42] proposed leader-follower attention, where audio, visual, and text modalities are combined to obtain the attended feature representations using the modality-wise attention scores, which is further concatenated to visual features for final predictions.
@@ 3.1 Visual Network
Facial expressions play a significant role in conveying the emotional state of a person. In videos, spatial information provide the semantic regions of face pertinent to the expressions, whereas temporal dynamics convey the evolution of expressions in videos across the frames. Therefore, effectively modeling the spatial and temporal dynamics of facial expressions in videos is crucial to obtain robust visual feature representations. Several approaches have been explored using 2D CNNs in conjunction with LSTMs, where 2D CNNs are used to encode the spatial information and LSTMs for temporal dynamics of facial expressions [45, 46]. With the advent of 3D CNN models [27], they are successfully explored for DER [12, 21, 22] and showed improvement over 2D CNNs with LSTMs. It has also been shown that 3DCNNs in combination with LSTMs are effective in capturing the temporal dynamics [23], where 3DCNNs are efficient in capturing short-term dynamics and LSTMs are robust in modeling long-term dynamics. Recently, TCNs are found to be promising to effectively capture the temporal dynamics for DER [42, 36]. In this work, we have used Resnet-50 [26] pretrained on MS-CELEB-1M dataset [47], which is further fine-tuned on FER+ [43] dataset similar to that of [42]. We have further used TCNs to effectively capture the temporal dynamics of facial expressions.
@@ 3.2 Audio Network
Speech-based emotion recognition is another promising research area due to the rich emotion-relevant information in the vocal expressions. With the advancement of deep learning models, vocal expressions are encoded using 1D CNNs with raw speech signals [10] or 2D CNNs with spectrograms [11, 12]. Some of the works also explored deep features in combination with conventional hand-crafted features to encode the vocal expressions [16, 48]. Recently, spectrograms have been widely explored as they are found to be efficient in capturing the affective states of the vocal expressions [42, 21]. Therefore, we have also explored spectrograms with 2D CNNs to encode the vocal expressions. Specifically, we have used VGG-Net architecture pretrained on large-scale audioset dataset [49]. Similar to visual modality, we have also used TCNs to encode the temporal dynamics of frame-level vocal embeddings.
@@ 3.3 Text Network
Text modality is another predominantly explored modality for emotion detection, which carries semantic emotion-relevant information in the text data [50]. Effectively leveraging the textual data can boost the performance of multimodal fusion as they can offer significant emotion-relevant information and complement audio and visual modalities. Based on transformers, BERT features are predominantly explored text encoders for emotion recognition in the literature [51]. Therefore, we also used BERT as text encoder, followed by TCNs to encode the temporal information across the word embeddings.
@@ 3.4 Recursive Joint Cross-Modal Attention
Given the input video sub-sequence $S$ of $K$ frames, the audio, visual and text data are preprocessed and fed to the corresponding encoders, followed by TCNs to obtain the feature representations of the respective modalities as shown in Fig 1. The feature representations of the audio, visual and text modalities are denoted by ${\boldsymbol{X}}_{a}\boldsymbol{=}\{\boldsymbol{x}_{a}^{1},\boldsymbol{x}_{a}^{2},...,\boldsymbol{x}_{a}^{K}\boldsymbol{\}}\in\mathbb{R}^{d_{a}\times K}$, ${\boldsymbol{X}}_{v}\boldsymbol{=}\{\boldsymbol{x}_{v}^{1},\boldsymbol{x}_{v}^{2},...,\boldsymbol{x}_{v}^{K}\boldsymbol{\}}\in\mathbb{R}^{d_{v}\times K}$ and ${\boldsymbol{X}}_{t}\boldsymbol{=}\{\boldsymbol{x}_{t}^{1},\boldsymbol{x}_{t}^{2},...,\boldsymbol{x}_{t}^{K}\boldsymbol{\}}\in\mathbb{R}^{d_{t}\times K}$ respectively, where $d_{a}$, $d_{v}$, and $d_{t}$ are dimensions of audio, visual and text features. $\boldsymbol{x}_{a}^{k}$, $\boldsymbol{x}_{v}^{k}$ and $\boldsymbol{x}_{t}^{k}$ represents the feature vectors of the individual frames of audio, visual, and text modalities.
@@ 4.1 Dataset
Affwild is the largest database in the field of affective computing, consisting of $298$ videos captured under extremely challenging conditions from YouTube videos [52]. The database is extended substantially to foster the development of robust models for estimating valence and arousal by introducing $260$ more videos, resulting in a total of $558$ videos with $1.4$M frames [53]. The database has been further expanded by adding new videos in the subsequent series of ABAW challenges [54, 55, 56, 57, 58, 59]. In the 6th ABAW challenge [60] for the valence-arousal track, the dataset is provided with $594$ videos of around $2,993,081$ frames from $584$ subjects. Sixteen of these videos display two subjects, both of which have been annotated. The final annotations are obtained by taking an average of annotations provided by four experts, using a joystick. The annotations for valence and arousal are provided continuously in the range of $[-1,1]$. The dataset is partitioned into the training, validation, and test sets in a subject-independent manner to ensure that each subject appears exclusively in one partition. The partitioning resulted in $356$, $76$, and $162$ videos for train, validation, and test partitions respectively.
@@ 4.2.1 Preprocessing
For the visual modality, we have used the cropped and aligned images provided by the challenge organizers [60]. For the missing faces in the video frames, we have considered black frames (i.e., zero pixels), and the video frames with no annotations of valence and arousal i.e., with annotations of $-5$ are excluded. The given video sequences are divided into sub-sequences of $300$ (i.e.,$K=300$) with a stride of $200$ frames and the facial images are resized to $48\times 48$.
@@
For the text modality, the extracted speech signals from audio preprocessing are fed to the pretrained speech recognition model of Vosk toolkit [^2] https://alphacephei.com/vosk/models/vosk-model-en-us-0.22.zip to obtain the recognized words along with word-level timestamps. Next, a pretrained punctuation restoration and capitalization model [^3] https://pypi.org/project/deepmultilingualpunctuation/ is used to restore the punctuations of the recognized words, which carries semantic information pertinent to emotional states. Now BERT features are extracted at word-level using a pre-trained BERT model [^4] https://pypi.org/project/pytorch-pretrained-bert/. The word-level features are computed by taking a summation of the last four layers of the BERT model [61]. The recognized words may usually span a larger time window of multiple frames. In order to synchronize the word-level BERT features of text modality with audio and visual modalities, the word-level text embedding is populated as per the timestamp of each word by reassigning the same word-level feature to all the frames within the time span of the corresponding word.